OpenStudy (anonymous):

Can someone walk me through how to solve this problem? (5+radical2) (5- radical 2)

OpenStudy (anonymous):

use the formula a^2 - b^2 = (a+b)(a-b)

OpenStudy (anonymous):

wait don't you foil?

OpenStudy (anonymous):

it will be a longer way

OpenStudy (anonymous):

can u show me how to foil it? Im confused on foiling with radicals so can u make sure I'm doing it correctly?(:

OpenStudy (anonymous):

sure, go ahead and solve it

OpenStudy (anonymous):

ok when i do \[5 \times -\sqrt{2}\]... im confused wut i get? -radical 10 or something else?

OpenStudy (anonymous):

u get -5(radical 2)

OpenStudy (anonymous):

\[-5\sqrt{2}\]

OpenStudy (anonymous):

and then also what do i get when I times radical 2 times -radical 2...is it -radical 2 or -radical 4

OpenStudy (anonymous):

\[\sqrt{2} * \sqrt{2} = 2\]

OpenStudy (anonymous):

so it would equal radical 4 and that equals 2 right?

OpenStudy (anonymous):

no it would equal -2

OpenStudy (anonymous):

oh i see... is there a certain rule i should know when multiplying radicals?

OpenStudy (anonymous):

since it is same as \[\sqrt{2}\times \sqrt{2}\times(-1) = -2\]

OpenStudy (anonymous):

wait so wut is the answer to this problem?

OpenStudy (anonymous):

25 - 2 = 23

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

can u help me with another problem?